OpenStudy (anonymous):

15. A car traveling at 50 m/s applies its brakes. The brakes slow the car at a rate of 3m/s2. If the car slows to a speed of 26 m/s, how far did it travel while the brakes were being applied?

OpenStudy (amistre64):

what physics formulas are you aware of for this?

OpenStudy (goformit100):

Welcome to Open Study .................. Sir/Ma'am @habibmiranda

OpenStudy (amistre64):

otherwise we can develop one using calculus

OpenStudy (anonymous):

Well, I know the basic x=.5at^2 and v=sqrt(2ax)

OpenStudy (anonymous):

I have not yet taken a calculus class... :(

OpenStudy (amistre64):

those are fine the x=.5at^2 will do fine, but needs to be modified at least in my mind.

OpenStudy (anonymous):

Ok. how so?

OpenStudy (amistre64):

a is causing the car to slow down, so do you agree that its going to be negative? and at a rate of 3?

OpenStudy (anonymous):

yea, I figured that.

OpenStudy (amistre64):

x=-3(.5) at^2 but we have a starting velocity, we start at 50t x = -3(.5) at^2 + 50t now, at t=0 we have a place to start to find the speed at any moment, i have to resort to calculus since its what i know best and can never recall the formulas lol v = -3(2)(.5)t + 50, for what value of t does v = 26?

OpenStudy (amistre64):

\[v-vi=at\] might be the formula?

OpenStudy (amistre64):

v_i that is

OpenStudy (amistre64):

starting to recall my physics class lol. \[v_f-v_i=at\]\[26-50=-3t\] there for\[x=-\frac{3}{2}t^2\] suffices for distance traveled.

OpenStudy (anonymous):

I guess I'm a little lost on how to find t.

OpenStudy (amistre64):

um ... divide by -3 of course.

OpenStudy (amistre64):

\[v_f=at+v_i\] \[v_f-v_i=at\] \[\frac{v_f-v_i}{a}=t\] if we simply wnat to sub this in then: \[x=\frac{a}{2}\left(\frac{v_f-v_i}{a}\right)^2\]

OpenStudy (anonymous):

that gives me a value of -8 for time and when I plug it in to find x it gives me 96 meters which is the wrong answer. The answer is supposed to be 304 meters

OpenStudy (amistre64):

26 - 50 = -24 -24/-3 = 8 ... but then 8^2 = -8^2 so that shouldnt be an issue x = .5(-3) * 8^2 x = -3(32) = 96 let me try it my calculus way ..... see if i get the same results.

OpenStudy (anonymous):

ok. sounds good. I appreciate all your help BTW.

OpenStudy (amistre64):

a = -3; integrate to get velocity ............................................... v = -3t + C such that v(0) = 50; C = 50 v = -3t + 50; integrate to get position when v=26, t=8 ................................................. s = -3/2 t^2 + 50t + K such that at s(0) we have a starting position of 0 to measure from. K = 0 s = -3/2 t^2 + 50t gives us our position when t=8 we get .... 304

OpenStudy (anonymous):

So would you say that in order to solve this problem we need to know calculus?

OpenStudy (amistre64):

need to know calculus? no. you just have to remember the proper formulas that the physics book give us ....

OpenStudy (amistre64):

i just work better from a calculus point of view :) since i cant remember a formula for the life of me lol

OpenStudy (anonymous):

I guess that's my problem. I don't recall what formula that would solve this from my book. Was it a derivation of a more common equation?

OpenStudy (amistre64):

one way to consider it is: a = -3 v = -3t; v = 50 when t=-50/3 v = 26 when t = -26/3 the position of x when t = -50/3 is: a the position of x when t = -26/3 is: b the difference between a and b is the distance traveled.

OpenStudy (anonymous):

Ok. That's actually very helpful.

OpenStudy (amistre64):

its an approach for using the stuff that we can recall :)

OpenStudy (anonymous):

Thanks again. You were awesome!

OpenStudy (amistre64):

youre welcome, and good luck :)